OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the hypotenuse in a right triangle whose legs are 3 and 4. A. 3 B. 5 C. 4 D. 2

OpenStudy (ivettef365):

that is the same as 3^2+4^2 = c^2, solve for c

OpenStudy (anonymous):

a^2 + b^2 = c^2 3^2 + 4^2 = c^2 Solve for c. 9 + 16 = c^2 25 = c^2 √25 = c 5 = C C = Hypotenuse Hypotenuse = 5

OpenStudy (anonymous):

5